Errors in measuring the time of arrival of a wave front from an acoustic source sometimes have an approximate beta distribution. Suppose that these errors, measured in microseconds, have approximately a beta distribution with α = 1 and β = 3.

What is the probability that the measurement error in a randomly selected instance is less than 0.8 µs?

Give the mean and standard deviation of the measurement errors in microseconds.
Answer #1

a)

f(x)=(1/B(α,β))*xα-1*(1-x)β-1 = 3*x^0*(1-x)^2 =3(1-x)2

P(X<0.8)= f(x )dx = 3(1-x)2 dx =-3*(1-x)3/3|0.80 =0.9920

b)

mean μ=E(X)=α/(α+β)=1/(1+3) = 0.2500
variance=σ2= αβ/((α+β)2*(α+β+1))= 0.0375
standard deviation=σ=√σ2= 0.1936